OpenStudy (anonymous):

Write the equation of the circle: center at origin passing at (2,4)

OpenStudy (noelgreco):

The first thing you want to do is draw a sketch so you can see how to calculate the radius of the circle.

OpenStudy (anonymous):

after drawing the circle what should i do next?

OpenStudy (anonymous):

do you know the equation of the circle: (x-a)^2 + (y-b)^2 = r^2?

OpenStudy (noelgreco):

Calculate the radius of the circle.

OpenStudy (anonymous):

I already computed for the radius of the circle but i don't know if it's right...the adius of my circle is 2 square root of 5???

OpenStudy (noelgreco):

That's correct. Now, since the center of the circle is at the origin, the standard form of the circle's equation is x^2 + y^2 = r^2

OpenStudy (anonymous):

My equation is x^2+y^2-20=0 in general form

OpenStudy (noelgreco):

That's correct. Usually the 20 would be on the right side, though.

OpenStudy (anonymous):

Ok. Thanks for the help :)